# Simple Hamiltonian dynamics is a powerful quantum processing resource

### v1

@@ Abstract
A quadrillion dimensional Hilbert space hosted by a quantum processor with over 50 physical qubits has been expected to be powerful enough to perform computational tasks ranging from simulations of many-body physics to complex financial modeling. Despite few examples and demonstrations, it is still not clear how we can utilize such a large Hilbert space as a computational resource; in particular, how a simple and small quantum system could solve non-trivial computational tasks. In this paper, we show a simple Ising model capable of performing such non-trivial computational tasks in a quantum neural network model. An Ising spin chain as small as ten qubits can solve a practical image classification task with high accuracy. To evaluate the mechanism of its computation, we examine how the symmetries of the Hamiltonian would affect its computational power. We show how the interplay between complexity and integrability/symmetries of the quantum system dictates the performance as quantum neural network.
@@ I Introduction
Recent developments in the field of quantum computation have seen processors composed of hundreds of qubits realized which can be used to simulate novel phenomena for a variety of physical, chemical and biological systems, as well solving interesting financial and mathematical problems [1, 2, 3, 4, 5, 6, 7]. These processors are large enough that they should be able to generate sufficiently complex quantum dynamics so that they can be used to simulate and explore the nature of quantum many-body physics [8, 9, 10, 11, 12, 13, 14]. If such an expectation were feasible, then we could consider whether it would be possible to reverse this process and use a typical many-body system, such as the Ising model, to perform computational tasks? This leads to a very interesting question: would natural quantum dynamics alone be a useful resource for quantum computation? If so, would the common beliefs for the requirements for quantum computation change, such as the necessity of non-integrable systems for quantum machine learning [15] and circular unitary ensemble (CUE) processes for the quantum reservoir [16, 17]? Do the commonly used measures for computational capacity, such as the information processing capacity [18, 19, 20] remain appropriate? In this article, we aim to answer these questions by designing a quantum computational model using only the Ising model for its quantum contribution.
@@
To explore and address these questions, we benchmark quantum neural networks using quantum extreme reservoir computation (QERC) [21]. This QERC process as depicted in Fig. 1 is an extreme machine learning model with a quantum reservoir for the middle hidden layer of nodes. As such the quantum reservoir is the only quantum component in the system. The information present in the quantum reservoir would be extracted and converted to the classical data during the measurement step, which is then fed to the one-layer classical neural network. The quantum reservoir provides a given quantum dynamics, while the total system works as the quantum neural network model with the dual nature associated with the extreme machine learning and reservoir computation. More specifically this is a two-layer neural network with a large node number and no optimization nor control on the recurrent neural network (the reservoir).
@@
Among the various quantum machine learning models, we choose the QERC for this benchmark for several reasons. First and foremost variational quantum algorithms [22, 23, 24], which are a class of quantum neural networks, have been extensibly studied for quantum phase transitions and determining quantum eigenstates [25, 26, 7]. However these models require the ability to apply a quantum circuit with a parameter which is optimized by classical computation. Furthermore there is a known hard optimization problem [27] to be solved to apply these models, besides the difficulty to implement a quantum circuit to realize a good feature space [28, 29]. Second, quantum reservoir computation [30, 31, 32, 33, 34, 35, 36] could be closer to the QERC model, however it was constructed as an quantum extension of classical reservoir computation and designed to analyze temporal data. Quantum extreme learning machine had also been conceptually proposed as an extension of the two-layer neural network to be replaced by a quantum neural network [23, 37]. Although, unlike the concept of the quantum perceptron [38, 39, 40], these models could be consistent with quantum mechanics, they are not be suitable as a benchmarking candidate until they show at least one concrete instance where they can demonstrate its capability to achieve a significantly high accuracy.
@@
Figure 1: Schematic illustration of the structure for quantum extreme reservoir computation (QERC). Here the blue and green areas represent the quantum and classical components, respectively. At the first layer, inputs (vectors) are transformed into vectors in a $2N$-dimensional space using principal component analysis (PCA). These components are then encoded into our quantum system using single qubit rotation gates. Quantum many-body dynamics (QMD), such as the Ising models, maps the input to complex final states. Probability distributions are then obtained at the measurement layer and sent to a single layer classical neural network. The training process is only applied to the classical neural network while the quantum reservoir implemented using QMD is fixed.
@@
By contrast, the QERC model has an explicit form of computation and requires only the initialization of the quantum reservoir to perform various tasks. The (not only conceptual but also) explicit form of the QERC is essential to evaluate the validity of the measures and requirements commonly used for quantum computational capacity. Moreover, it has numerically demonstrated the accuracy necessary for benchmarking with few qubits. This is essential for the numerical evaluation. Now the estimation of the quantum computational power of quantum machine learning models tends to have its focus on the complexity of the quantum process. However, a quantum neural network model provides its capability through the process from the classical input to the classical output. Hence the complexity of the quantum process is not enough to evaluate the capability of the quantum neural network. In the field of computer science benchmarking has long been used for the evaluation of computational capability.
@@
In this paper, we employ benchmarking of the quantum reservoirs within the QERC model to enable us to evaluate the requirements and capacity measures of the quantum reservoir. To achieve this, we must first fix the nature of the output measurements. As such measurements can be as computationally powerful operations, similar to the universality of gate-based quantum computation [41], our measurements in this QERC model will be a projective measurements on to the computational basis. These can be implemented using single qubit $Z$-axis measurements. Next, the system size in our benchmarking will be around 10 qubits. As such we need to choose an appropriate task to run our benchmarking on, and the MNIST data set for hand-written numbers [42] has an sufficient complexity for this size regime.
@@ II Quantum reservoir models
It is well known that any quantum operation can be decomposed with an arbitrary accuracy with a universal gate set involving only single and two qubit operations [43, 44]. It has been proven that multi-qubit operations are not necessary [43, 44]. However, in Hamiltonian based computation, there is a complexity difference between two-body interactions and those above. In previous work [21], a discrete time crystal model was used corresponding to a multi-body Hamiltonian [21]. It is not therefore clear that two-body Hamiltonian are enough to achieve the same performance in QERC. To evaluate this, let us first employ two Ising models that consist only of one- and two-qubit interactions: the ZZ-Ising model with the corresponding Hamiltonian,
@@
where $\sigma_{l}^{\mu}$ $(\mu=x,y,z)$ are the usual spin-1/2 Pauli matrices at the $l$-th site. Next $J_{lm}$ and $g\geq 0$ are the strength of the two-qubit interaction and the transverse magnetic field, respectively. We set as $J_{lm}=J_{0}/|l-m|^{\alpha}$, where $J_{0}$ is the strength of the interaction and $\alpha$ a long range order parameter. These Hamiltonians have been experimentally demonstrated in a variety of different physical platforms including [45, 8, 46, 47, 48, 49, 50, 51]. In those trapped ions demonstrations  [45, 47, 48] $\alpha\sim 1.5$.
Now the XX-Ising and ZZ-Ising models are related by the unitary transformation $\hat{H}_{XX}=\exp{(i\sum_{l}\frac{\pi}{4}\sigma_{l}^{y})}\hat{H}_{ZZ}\exp{(-i\sum_{l}\frac{\pi}{4}\sigma_{l}^{y})}$. However unlike the situation for the ZZ-interaction, the XX-interaction breaks the U(1)-symmetry meaning the properties and capacity as a quantum reservoir may be different as we observe later.
Figure 2: Depiction of quantum computational circuit for the ZZ-X and Cliffort+T models. In (a) for the ZZ-X model, the ZZ-interaction part can be decomposed to a circuit of quantum gates, however it can be advantageous for the ZZ-X model to have the ZZ-interaction part created using Hamiltonian dynamics which would not require any two-qubit gates. As such the simplicity of the model allows one to use more specialized quantum computational devices such as quantum annealer. Next (b) depicts the Cliffort+T model where the Clifford circuit consists of only of randomly selected Clifford gates, followed by one layer of T-gates before measurement.
Next given the XX-Ising and ZZ-Ising models let us consider two further models. The first is the ZZ-X model who dynamics are governed by
$\hat{U}_{\mathrm{ZZX}}=\exp\left(-i\theta_{x}\sum_{l}\sigma_{l}^{x}\right)\exp\left(-i\sum_{l<m}\theta^{J}_{lm}\sigma_{l}^{z}\sigma_{m}^{z}\right)$ (3)
where $\theta_{l}^{x}$ and $\theta_{lm}^{J}$ are rotation angles. To simplify our analysis further we set $\theta_{l}^{x}=\theta_{x}$ to be equal over the site, similar to the transverse field of the Ising model. Now when $\pi/2\leq\theta_{x}<\pi$, the rotation part can be written as,
@@
where $\hat{P}_{x}$ is the parity operator. Because $\hat{P}_{x}$ does not affect the computational basis set (it is a bijection between computational basis sets), the performance of $\pi/2\leq\theta_{x}<\pi$ will be the same as that of $20\leq\theta_{x}-{\pi}/{2}<\pi/$ 2. Thus, in the analysis, we investigate only the $20\leq\theta_{x}\leq\pi/$ 2 range. Next the angle $\theta_{lm}^{Jt}$ is site-dependent and like for the Ising model, we consider the long-range interaction to be represented by $\theta^{J}_{lm}=\theta_{J}/|l-m|^{\alpha}$ with $\theta_{J}$ being the maximum angle.
@@
To be more computational complexity oriented, our second additional model is associated with a quantum circuit of the random Clifford gate circuit plus one depth of T-gates for each qubit as the quantum reservoir. Fig. 2 depicts our various strategies where the ZZ-X model has two circuits: the natural dynamics of ZZ-Ising model and the single X rotation on each qubit a seen in Fig. 2 (a). Next Fig. 2 (b) shows the fully gate-decomposed model with random Clifford circuit and the T-gate layer. This is referred to as the Random Cilifford+T model in the remainder of this paper.
@@
Finally we employ Haar measure sampling of the unitary matrices as the standard model to evaluate others against. To evaluate this non-trivial dynamics of quantum computation, randomness is often used as an estimator for the computational capability in both experimental demonstrations [52, 13] and quantum variational algorithms [29, 53]. Haar measure sampling has also been used to design quantum machine learning models [16] which are deeply connected to information scrambling in the computational process. The complication in the evaluation here is that the quantum reservoir is not the only element to process the task, that is, the role of the quantum reservoir can change dependent upon the nature of the quantum reservoir. Haar measure random sampling may be ideal for information scrambling, however it is not necessary easy to extract information from such a final state. In quantum algorithms with single-shot projective measurement, information on the quantum computer needs to be scrambled during the process, however equally needs to be converged to the solution of the problem. In QERC, and more generally any quantum machine learning model, the role of each component is not clear. For instance, usually the circuit before the variational model in VQA is considered for encoding and information scrambling, however that is only a convenient recognition to design the model. Similarly, in QERC it is convenient to think that the quantum reservoir is their to exploit the large Hilbert space indicating information scrambling, however there is no necessity for the model to divide the task in this way. Hence, it is necessary to include Haar measure random sampling in the benchmark to understand the mechanism of the computation.
@@ III Benchmarking
In this benchmarking we will evaluate the quantum reservoir models described in the previous section as they are used in the QERC shown in Fig. 1. It is important to remember that in the QERC model, there are two extra network layers: the encoding of classical data to the quantum reservoir and the measurement of the quantum reservoir. As we use the MNIST hand-written digits (0 to 9) data set for our benchmark, we need to use the principal component analysis (PCA) to encode the $2N$ largest principal components to $N$ spins by single rotations on the initialized state $|0\rangle\otimes\cdots\otimes|0\rangle$ as
@@
with $0\leq\theta_{l},\phi_{l}\leq\pi$ [21]. For each image, the first $N$ elements are encoded on the angles $\theta_{l}$, while the next $N$ components are encoded on the phase $\phi_{l}$. The smaller components are ignored. In this paper, we consider $10$ sites for the Ising model with $10$ spins as qubits.
The quantum reservoir readout is achieved by single-qubit projective measurements on the computational basis at the M-layer [21]. The output classical data is then treated and processed by the one-layer classical neural network performing the gradient decent AdaGrad [54] machine learning algorithm as shown in Fig 1 (see appendix A for details on the classical neural network). It is important to note that the M-layer is a middle layer in the neural network of the QERC, and the measurement takes a computational role. In this paper, the computational basis states 0 and 1 are the eigenstates of the $Z$ operator for each qubit.
@@ IV Performance
Let us begin takeing the ZZ-Ising model as the quantum reservoir of the QERC to evaluate its performance. Fig. 3 shows the results for the two different strategies with this quantum reservoir. The light blue dots (training) and the light green dots (testing) are for the ZZ Ising model while the blue dots (training) and the green dots (testing) are for the ZZ Ising model with the randomization of single qubit rotation before the time-evolution of the quantum reservoir as shown in the inset (b-2). We have allowed the system to evolve for a a scaled time $J_{0}t=3.5$.
@@
Figure 3: Plot of the $g$-dependence of the ZZ-Ising model for a scaled time $J_{0}t=3.5$ in (a) with ten sites, where the transverse magnetic field is applied along the x-axis. The blue and green lines show the training and testing results for the encoder shown in the inset (b-1), while the cyan and light-green lines show the training and testing results for the encoder with random single qubit rotation shown in (b-2).
@@
Although the training accuracy for the Haar measure sampling (indicated as the solid red line in Fig. 3 is slightly better than the highest accuracy with the ZZ Ising model, the highest accuracy with the ZZ-Ising model at $g/J_{0}\sim 1$ achieves the accuracy as high as the testing accuracy given by Haar measure random sampling. The degradation of accuracy in the strong transversal magnetic field regime $(g/J_{0}>1)$ can be overcome by the randomization of the single qubit rotation. It is worth noting that the single qubit gates alone (without the quantum reservoir) reduces the accuracy rate in this model. This indicates that this process does not directly contribute to the computation, but converts the input state in a form that the quantum reservoir can be more effective. Although the transversal magnetic field is necessary for the QERC to perform, the strong transversal magnetic field introduces the bias in the quantum dynamics, which may cause a mismatch between the tendency in the input data and the quantum reservoir similar to the behavior observed in classical machine learning  [55]. Another important aspect of the performance is that the ZZ-Ising model has a superior generalization performance around $g/J_{0}\sim 1$.
@@
Next we explore the QERC’s performance using the XX-Ising model given in Eq.  2. We begin by showing in Fig. 4 (a) the g-dependence of the XX-Ising model, where we clearly observe that its performance is high even in the $g/J_{0}<1$ regime. That shows that the XX-Ising model can utilize the many-body interaction without the transverse field for computation. The effect of the transverse field appears when we analyze the $\alpha$-dependence of the performance. The accuracy rates slightly drop for larger $g/J_{0}$ when $\alpha$ is large. The best accuracy for large $\alpha$ occurs for $g/J_{0}=0.5$ as shown in Fig. 4 (b).
Figure 4: Plot of the $g$ and $\alpha$-dependence of the XX-Ising models for a scaled time $J_{0}t=3.5$. Here (a) illustrates the result of the g-dependence with $\alpha=1.5$ while (b) shows the $\alpha$-dependence with $g/J_{0}=0.5$. The blue and green dashed lines with dots plot the training and testing accuracy with $\alpha$ ranging from $\alpha=0.0$ to $10$ (the blue and green solid lines indicate the training and testing accuracy with $\alpha=\infty$).
@@
Now, let us explore our two models that have a quantum circuit structure. Our first evaluation is for the ZZ-X model given by Eq. 3, noting that this model is the exactly the same as the ZZ-Ising model without the transversal field. It is interesting to see if the single qubit rotation is enough to extract the potential complexity generated through the ZZ-Ising interaction. Fig. 5 (a) illustrates the performance of the ZZ-X model dependency on the angle $\theta_{x}$ for each different angle $\theta_{J}$, while (b) shows the $\alpha$ dependency on the performance. Next (c) shows the performance dependency on the system size for $\theta_{J}=2\pi$. The wide range of $\theta_{x},\alpha$ allows us to achieve high accuracy performance for our classification task.
model Train acc. (std.) Test acc.
@@
T (HTH) $0.867\pm 0.001$ $0.8627\pm 0.001$
Clifford $0.902\pm 0.001$ $0.884\pm 0.002$
Clifford+T $0.990\pm 0.000$ $0.970\pm 0.001$
Table 1: Training and testing reservoir accuracy rates for models.
Our second quantum circuit structure model is the Clifford+T model and we present its performance in Table 1 along with the other models we have considered. We clearly observe that the Clifford+T model has the best performance in terms of both the training and testing accuracy. The ZZ-Ising, XX-Ising and ZZ-X models are however within 0.5%.

### v2

# Simple Hamiltonian dynamics is a powerful resource for image classification

@@ Abstract
A quadrillion-dimensional Hilbert space hosted by a quantum processor with over 50 physical qubits has been expected to be powerful enough to perform computational tasks ranging from simulations of many-body physics to complex financial modeling. Despite a few examples and demonstrations, it is still unclear how we can utilize such a large Hilbert space as a computational resource, particularly how a simple and small quantum system could solve non-trivial computational tasks. This paper shows a simple Ising model capable of performing such non-trivial computational tasks in a quantum neural network model. An Ising spin chain as small as ten qubits can solve a practical image classification task with high accuracy. To evaluate the mechanism of its computation, we examine how the symmetries of the Hamiltonian would affect its computational power. We show how the interplay between complexity and symmetries of the quantum system dictates the performance as quantum neural network.
@@ I Introduction
Recent developments in the field of quantum computation have seen processors composed of hundreds of qubits realized which can be used to simulate novel phenomena for a variety of physical, chemical, and biological systems, as well as solving interesting financial and mathematical problems [1, 2, 3, 4, 5, 6, 7]. These processors are large enough that they should be able to generate sufficiently complex quantum dynamics so that they can be used to simulate and explore the nature of quantum many-body physics [8, 9, 10, 11, 12, 13, 14]. If such an expectation were feasible, then we could consider whether it would be possible to reverse this process and use a typical many-body system, such as the Ising model, to perform computational tasks? This leads to a very interesting question: would natural quantum dynamics alone be a useful resource for quantum computation? If so, would the common beliefs for the requirements for quantum computation change, such as the necessity of non-integrable systems for quantum machine learning [15] and circular unitary ensemble (CUE) processes for the quantum reservoir [16, 17]? Do the commonly used measures for computational capacity, such as the information processing capacity [18, 19, 20] remain appropriate? In this article, we aim to answer these questions by designing a quantum computational model using only the Ising model for its quantum contribution.
@@
To explore and address these questions, we benchmark quantum neural networks using quantum extreme reservoir computation (QERC) [21]. This QERC process, as depicted in Fig. 1, is an extreme machine learning model with a quantum reservoir for the middle hidden layer of nodes. As such, the quantum reservoir is the only quantum component in the system. The information in the quantum reservoir would be extracted and converted to classical data during the measurement step. It is then fed to a one-layer classical neural network, a linear classifier for classification problems. The quantum reservoir provides a given quantum dynamics, while the total system works as a quantum neural network model with the dual nature associated with extreme machine learning and reservoir computation. More specifically, this is a two-layer neural network with a large node number and no optimization nor control on the recurrent neural network (the reservoir).
@@
Among the various quantum machine learning models, we choose the QERC for this benchmark for several reasons. First and foremost, variational quantum algorithms [22, 23, 24], which are a class of quantum neural networks, have been extensively studied for quantum phase transitions and determining quantum eigenstates [25, 26, 7]. However, these models require the ability to design a quantum circuit with a parameter that is optimized by classical computation. That optimization is a known hard optimization problem in these models [27] . Further, it is generally difficult to implement the quantum circuit needed to realize the required feature space [28, 29]. Second, quantum reservoir computation [30, 31, 32, 33, 34, 35, 36] could be closer to the QERC model. However, it was constructed as a quantum extension of classical reservoir computation and designed to analyze temporal data. Quantum extreme learning machine had also been conceptually proposed as an extension of the two-layer neural network to be replaced by a quantum neural network [23, 37]. Although, unlike the concept of the quantum perceptron [38, 39, 40], these models could be consistent with quantum mechanics, they are not suitable as a benchmarking candidate until they show at least one concrete instance where they can demonstrate their capability to achieve a significantly high accuracy.
@@
Figure 1: Schematic illustration of the structure for quantum extreme reservoir computation (QERC). Here, the blue and green areas represent the quantum and classical components, respectively. At the first layer, inputs (vectors) are transformed into vectors in a $2N$-dimensional space using principal component analysis (PCA). These components are then encoded into our quantum system using single-qubit rotation gates. Quantum many-body dynamics (QMD), such as the Ising models, map the input to complex final states. The measurement obtains probability distributions and is then transformed to a distribution with zero mean and one standard deviation. We refer to the measurement and this transfer together as the M-layer. Then, the distribution is sent to a linear classifier. The training process is only applied to the classical neural network, while the quantum reservoir implemented using QMD is fixed.
@@
By contrast, the QERC model has an explicit form of computation and requires only the initialization of the quantum reservoir to perform various tasks. The (not only conceptual but also) explicit form of the QERC is essential to evaluate the validity of the measures and requirements commonly used for quantum computational capacity. Moreover, it has numerically demonstrated the accuracy necessary for benchmarking with few qubits. This is essential for the numerical evaluation. Now, the estimation of the quantum computational power of quantum machine learning models tends to focus on the complexity of the quantum process. However, a quantum neural network model provides its capability through the process from the classical input to the classical output. Hence, the complexity of the quantum process is not enough to evaluate the capability of the quantum neural network. In the field of computer science, benchmarking has long been used for the evaluation of computational capability.
@@
In this paper, we employ benchmarking of the quantum reservoirs within the QERC model to evaluate their requirements and capacity measures. To achieve this, we must first fix the nature of the output measurements. As such measurements can be as computationally powerful operations, similar to the universality of gate-based quantum computation [41], our measurements in this QERC model will be a projective measurements on to the computational basis. These can be implemented using single-qubit $Z$-axis measurements. Next, the system size in our benchmarking will be around 10 qubits. As such, we need to choose an appropriate task to run our benchmarking on, and the MNIST data set for hand-written numbers [42] has sufficient complexity for this size regime.
@@ II Quantum reservoir models
It is well known that any quantum operation can be decomposed with an arbitrary accuracy with a universal gate set involving only single and two-qubit operations [43, 44]. It has been proven that multi-qubit operations are unnecessary [43, 44]. However, in Hamiltonian-based computation, its complexity depends on the types of many-body interactions involved [45, 46]. For example, a system with classical many-body interactions and a transverse magnetic field, such as the Ising model used in adiabatic quantum computation, shows a different complexity between the two-body situation and those above [47, 46]. The previous work [21] used a discrete time crystal model consisting of a time-periodic Hamiltonian with only two-body interactions, but its effective Hamiltonian has more than three-body interactions. It is not, therefore, clear that a two-body Hamiltonian is enough to achieve the same performance in QERC. To evaluate whether two-body Hamiltonians are sufficient, let us first employ two Ising models that consist only of one- and two-body interactions: the ZZ-Ising model with the corresponding Hamiltonian,
@@
where $\sigma_{l}^{\mu}$ $(\mu=x,y,z)$ are the usual spin-1/2 Pauli matrices at the $l$-th site. Next $J_{lm}$ and $g\geq 0$ are the strength of the two-qubit interaction and the transverse magnetic field, respectively. We set $J_{lm}=J_{0}/|l-m|^{\alpha}$, where $J_{0}$ is the strength of the interaction and $\alpha$ is a long-range order parameter. These Hamiltonians have been experimentally demonstrated in a variety of different physical platforms, including [48, 8, 49, 50, 51, 52, 53, 54]. In those trapped ions demonstrations  [48, 50, 51] $\alpha\sim 1.5$..
Now the XX-Ising and ZZ-Ising models are related by the unitary transformation $\hat{H}_{XX}=\exp{(i\sum_{l}\frac{\pi}{4}\sigma_{l}^{y})}\hat{H}_{ZZ}\exp{(-i\sum_{l}\frac{\pi}{4}\sigma_{l}^{y})}$. Both Hamiltonians have a U(1) symmetry; however, only the ZZ-interaction shares the symmetry with the computational basis measurement. This difference can affect the performance of a quantum reservoir, as we will observe later.
Figure 2: Depiction of the quantum computational circuit for the ZZ-X and Cliffort+T models. In (a) for the ZZ-X model, the ZZ-interaction part can be decomposed into a circuit of quantum gates. However, it can be advantageous for the ZZ-X model to have the ZZ-interaction part created using Hamiltonian dynamics, which would not require any gate-circuit decomposition. As such, the model’s simplicity allows one to use more specialized quantum computational devices such as quantum annealer. Next (b) depicts the Clifford+T model, where the Clifford circuit consists of only randomly selected Clifford gates, followed by one layer of T-gates before measurement.
Next, given the XX-Ising and ZZ-Ising models, let us consider two further models. The first is the ZZ-X model whose dynamics are governed by
$\hat{U}_{\mathrm{ZZX}}=\exp\left(-i\sum_{l}\theta^{x}_{l}\sigma_{l}^{x}\right)\exp\left(-i\sum_{l<m}\theta^{J}_{lm}\sigma_{l}^{z}\sigma_{m}^{z}\right)$ (3)
where $\theta_{l}^{x}$ and $\theta_{lm}^{J}$ are rotation angles. To simplify our analysis further, we set $\theta_{l}^{x}=\theta_{x}$ to be equal over the site, similar to the transverse field of the Ising model. Now when $\pi/2\leq\theta_{x}<\pi$, the rotation part can be written as,
@@
where $\hat{P}_{x}$ is the parity operator. Because $\hat{P}_{x}$ is a bit-flip operation and the computational basis measurement is followed, the permutation can be absorbed into a linear transformation within the linear classifier. Thus, we investigate only the $20\leq\theta_{x}\leq\pi/$ 2 range in the analysis. Next, the angle $\theta_{lm}^{Jt}$ is site-dependent, and like for the Ising model, we consider the long-range interaction to be represented by $\theta^{J}_{lm}=\theta_{J}/|l-m|^{\alpha}$ with $\theta_{J}$ being the maximum angle.
@@
To be more computational complexity oriented, our second additional model is associated with a quantum circuit of the random Clifford gate circuit plus one depth of T-gates for each qubit as the quantum reservoir. Fig. 2 depicts our various strategies where the ZZ-X model has two circuits: the natural dynamics of the ZZ-Ising model and the single X rotation on each qubit as seen in Fig. 2 (a). Next, Fig. 2 (b) shows the fully gate-decomposed model with a random Clifford circuit and the T-gate layer. This is referred to as the Random Cilifford+T model in the remainder of this paper.
Finally, we employ Haar measure sampling of the unitary matrices as the standard model to evaluate others against. To evaluate this non-trivial dynamics of quantum computation, randomness is often used as an estimator for the computational capability in both experimental demonstrations [55, 13] and quantum variational algorithms [29, 56]. Haar measure sampling has also been used to design quantum machine learning models [16], which are deeply connected to information scrambling in the computational process.
@@
Haar measure random sampling may be ideal for information scrambling. However, it is not necessarily easy to extract information from such a final state. In the process of quantum computation, we expect to utilize the Hilbert space to gain quantum advantage, and the process for the input state to be spread in the Hilbert space is called information scrambling. Then at the stage of the readout, ideally the solution state is to be one of the computational basis, which is then read out by a single-shot computational basis measurement. Hence, in general we have two factors in quantum computation: information scrambling and conversion to/towards the solution. The complication of quantum-classical hybrid models is that we do not know which part of the model carries what role, or even if we define each role to each part of the model. Generally, quantum machine learning models including QERC face this complication. For instance, usually, the circuit before the variational model in VQA is considered for encoding and information scrambling, however, that is only a convenient recognition to design the model. Similarly, in QERC, it is convenient to think that the quantum reservoir is there to exploit the large Hilbert space, indicating information scrambling. However, there is no necessity for the model to divide the task in this way. Hence, it is necessary to include Haar measure random sampling in the benchmark to understand the mechanism of the computation. We choose a particular Haar-random unitary in this case.
@@ III Benchmarking
In this benchmarking, we will evaluate the quantum reservoir models described in the previous section as they are used in the QERC shown in Fig. 1. It is important to remember that in the QERC model, there are two extra network layers: the encoding of classical data to the quantum reservoir and the measurement of the quantum reservoir. As we use the MNIST hand-written digits (0 to 9) data set for our benchmark, we need to use the principal component analysis (PCA) to encode the $2N$ largest principal components to $N$ spins by single rotations on the initialized state $|0\rangle\otimes\cdots\otimes|0\rangle$ as
@@
with $0\leq\theta_{l},\phi_{l}\leq\pi$ [21]. For each image, the first $N$ elements are encoded on the angles $\theta_{l}$, while the next $N$ components are encoded on the phase $\phi_{l}$. The smaller components are ignored. This paper considers $10$ sites for the Ising model with $10$ spins as qubits.
The quantum reservoir readout is achieved by single-qubit projective measurements on the computational basis at the M-layer [21]. The output classical data is then treated and processed by the one-layer classical neural network performing the gradient descent AdaGrad [57] machine learning algorithm as shown in Fig 1 (see appendix A for details on the classical neural network). It is important to note that the M-layer is a middle layer in the neural network of the QERC, and the measurement takes a computational role. In this paper, the computational basis states 0 and 1 are the eigenstates of the $Z$ operator for each qubit.
@@ IV Performance
Let us begin by taking the ZZ-Ising model as the quantum reservoir of the QERC to evaluate its performance. Fig. 3 shows the results for the two different strategies with this quantum reservoir. The light blue dots (training) and the light green dots (testing) are for the ZZ Ising model while the blue dots (training) and the green dots (testing) are for the ZZ Ising model with the randomization of single qubit rotation before the time-evolution of the quantum reservoir as shown in the inset (b-2). We have allowed the system to evolve for a scaled time $J_{0}t=3.5$.
@@
Figure 3: (Plot of the $g$-dependence of the ZZ-Ising/XX-Ising models for a scaled time $J_{0}t=3.5$. (a) shows the $g$-dependence of the ZZ-Ising model with ten sites, where the transverse magnetic field is applied along the x-axis. The blue and green lines show the training and testing results for the encoder in the inset (b-1), while the cyan and light-green lines show the training and testing results for the encoder with random single qubit rotation shown in (b-2). (c) illustrates the result of the g-dependence with $\alpha=1.5$ of the XX-Ising model. The points and shadows are the mean value and the standard deviation taken from 40 to 50 epochs to reveal the stability of the optimization.
@@
Although the training accuracy for the Haar measure sampling (indicated as the solid red line in Fig. 3) is slightly better than the highest accuracy with the ZZ Ising model, the highest accuracy with the ZZ-Ising model at $g/J_{0}\sim 1$ achieves the accuracy as high as the testing accuracy given by Haar measure random sampling. The degradation of accuracy in the strong transversal magnetic field regime $(g/J_{0}>1)$ can be overcome by the randomization of the single qubit rotation. It is worth noting that the single-qubit gates alone (without the quantum reservoir) reduce the accuracy rate in this model. This indicates that this process does not directly contribute to the computation but converts the input state in a form that the quantum reservoir can be more effective. Although the transversal magnetic field is necessary for the QERC to perform, the strong transversal magnetic field introduces the bias in the quantum dynamics, which may cause a mismatch between the tendency in the input data and the quantum reservoir similar to the behavior observed in classical machine learning  [58].
@@
Next, we explore the QERC’s performance using the XX-Ising model given in Eq.  2. We begin by showing in Fig. 3 (c) the g-dependence of the XX-Ising model, where we observe that its performance is high even in the $g/J_{0}<1$ regime. That shows that the XX-Ising model can utilize the many-body interaction without the transverse field for computation.
This result indicates that only the two-body interaction is enough to generate a powerful reservoir with high explanation power. However, since ZZ-Interaction has U(1)-symmetry along the z-axis, which is the same as our measurement axis, its power does not appear when $g=0$ in the ZZ-Ising mode. In summary, in QERC, when generating strong quantum reservoirs from quantum dynamics, complex interactions and manipulations such as three-body and driven interactions are essentially unnecessary, and two-body interactions are sufficient. While, as in the case of ZZ-Ising, the system’s symmetry may mask the interactions’ ability, and additional manipulations or interactions are needed to eliminate or reduce the effect of symmetry.
@@
Now, let us explore our two models that have a quantum circuit structure. Our first evaluation is for the ZZ-X model given by Eq. 3, noting that this model is exactly the same as the ZZ-Ising model without the transversal field. It is interesting to see if the single qubit rotation is enough to extract the potential complexity generated through the ZZ-Ising interaction. Fig. 4 (a) illustrates the performance of the ZZ-X model dependency on the angle $\theta_{x}$ for each different angle $\theta_{J}$, while (b) shows the $\alpha$ dependency on the performance. Next (c) shows the performance dependency on the system size for $\theta_{J}=2\pi$. The dip in the accuracy rate around $\theta_{x}=\pi/4$ in Fig. 4 (a) for $N=10$ is smaller for a larger number of qubits. As the number of qubits increases, the wider range of $\theta_{x},\alpha$ allows us to achieve a high accuracy performance for our classification task.
model Train acc. (std.) Test acc. (std.)
Haar $0.990\pm 0.0004$ $0.969\pm 0.0004$
SU(2) random circuit $0.812\pm 0.0011$ $0.811\pm 0.0015$
@@
T gates (HTH) $0.867\pm 0.001$ $0.8627\pm 0.001$
Random Clifford $0.902\pm 0.001$ $0.884\pm 0.002$
Random Clifford+T $0.990\pm 0.000$ $0.970\pm 0.001$
Linear Classifier $0.936\pm 0.00029$ $0.928\pm 0.0007$
Linear Classifier (20 pc) $0.874\pm 0.00025$ $0.879\pm 0.0007$
Table 1: Training and testing reservoir accuracy rates for models. Here, as the linear classifier, we used a single-perceptron with the softmax activation function. The last one is the result of the linear classifier with PCA (20 pc).
Our second quantum circuit structure model is the Clifford+T model and we present its performance in Table 1 along with the other models we have considered. We clearly observe that the Clifford+T model has the best performance in terms of both the training and testing accuracy. The ZZ-Ising, XX-Ising, and ZZ-X models are, however, within 0.5%.